Two in-phase loudspeakers, which emit sound in all directions, are sitting side by side. One of them is moved sideways by 6.0 m, then forward by 6.0 m. Afterward, constructive interference is observed 1/4 and 3/4 of the distance between the speakers along the line that joins them.

Part A

What is the maximum possible wavelength of the sound waves?
Answer #1

We needto calculate the distance between the speakers,

d = sqrt(6^2 + 6^2) = 8.485 m

at 1/4 of distance between the speakers, Delta x = 3d/4 - d/4 = d/2

for constructive interference

Delta x = n x lambda

d/2 = n x lambda

8.485/2 = n x lambda   

for n = 1

lambda = 4.24 m

Comment in case any doubt please rate my answer.....